Circular No. 8444

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004gf, 2004gg, 2004gh, 2004gi Four new apparent supernovae have been reported from unfiltered

CCD images, 2004gf by T. Boles (cf. IAUC 8437), and the rest by M. Moore, K. Shimasaki, and W. Li (LOSS/KAIT; cf. IAUC 8443).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004gf Nov. 20.872 21^{h}57^{m}52.^{s}39 +42^{o}18′57′′.8 17.2 26′′.3 W, 35′′.6 N
2004gg Nov. 16.50 9 46 49.10 +16 02 44.5 18.5 6′′.1 W, 10′′.7 N
2004gh Nov. 16.55 10 24 31.55 −23 33 18.2 17.6 0′′.7 E, 7′′.5 S
2004gi Nov. 17.55 10 39 23.34 −27 54 40.0 14.9 21′′.4 W, 4′′.8 N

Additional magnitudes from the respective discoverers: SN 2004gf in UGC 11864, 1989 July 9, [20.0 (blue, Digitized Sky Survey); Sept. 3, [20.5 (red, Digitized Sky Survey); 2004 Oct. 4, [19.5; Nov. 21.871, 17.2. SN 2004gg in UGC 5234, May 2.21, [19.5; Nov. 22.58, 18.7. SN 2004gh in MCG –04-25-6, May 4.20, [19.0; Nov. 22.56, 17.6. SN 2004gi in MCG –05-25-32, May 4.21, [19.5, Nov. 22.58, 15.1. SN 2003W also appeared in UGC 5234, somewhat closer to the galaxy’s center.

COMET C/2004 RG_{113} (LINEAR) M. Bezpalko, Lincoln Laboratory, reports the LINEAR discovery of

an apparent comet with a short tail in p.a. ∼ 195^{o} on Nov. 20.435 UT. This object was identified by the Minor Planet Center with the apparently asteroidal object 2004 RG_{113} (also found by LINEAR; cf. MPS 113143 — Sept. 19 batch), with the discovery observations given below.

2004 UT  α_{2000}  δ_{2000}  Mag.
Sept. 6.39770  5^{h}30^{m}15.^{s}54  +41^{o}08′17′′.9  19.2
Nov. 20.43521  7 32 14.63  +53 40 09.8  18.4

The available observations, the following orbital elements, and an ephemeris appear on MPEC 2004-W38.

Epoch = 2005 Mar. 11.0 TT

T = 2005 Mar. 3.7235 TT  ω = 125.^{o}3104
e = 0.997212  Ω = 8.7720 }2000.0
q = 1.942281 AU  i = 21.6180
2004 November 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green